Question 19, chap 124, sect 7. part 1 of 4 10 points A solid sphere of radius 29 cm has a to- tal positive charge of 28.3 /C uniformly dis- tributed throughout its volume. Calculate the magnitude of the electric field at the center of the sphere.Answer in units ofN/C uestion 20, chap 124, sect 7. part 2 of 410 points Calculate the magnitude of the electric field 7.25 cm from the center of the sphere Answer in units of N/C Question 21, chap 124, sect 7. part 3 of 410 points Calculate the magnitude of the electric field 29 c Answer in units ofN/C n from the center of the sphere Question 22, chap 124, sect 7. part 4 of 410 points Calculate the magnitude of the electric field 57.7 cm from Answer in units of N/C the center of the sphere
Answer #3

.24ず E= 3.028x10% Co-294B (n299) 6 .) 1:5부.xcm: 0.5** 0,54キづ
Answer #2

Electric field in and out of the sphere when charges is uniformly distributed throughout it's volume

ours 尺 0.21) E- (98,09 ) с 28, 3x16-63 0.29) 6.577)
Answer #1

Solution 19)

At the center of the sphere, acc to Gauss law

Electric field, E=0

=======

Solution 20)

r = 7.25 cm = 0.0725 m

We know,

E(r) = Q/4 pi E0 r^2  

E(0.0725) = 9 x 10^9 x 28.3 x 10-6 / (0.0725)^2 = 2.1 x 106 N/C

Hence, E=4.84*10^7 N/C (Ans)

======

Acc to HOMEWORKLIB RULES, I can answer these many questions at a time only. Good luck!:)